### v1

# Diorama: Unleashing Zero-shot Single-view 3D Scene Modeling

Abstract
Reconstructing structured 3D scenes from RGB images using CAD objects unlocks efficient and compact scene representations that maintain compositionality and interactability. Existing works propose training-heavy methods relying on either expensive yet inaccurate real-world annotations or controllable yet monotonous synthetic data that do not generalize well to unseen objects or domains. We present Diorama, the first zero-shot open-world system that holistically models 3D scenes from single-view RGB observations without requiring end-to-end training or human annotations. We show the feasibility of our approach by decomposing the problem into subtasks and introduce robust, generalizable solutions to each: architecture reconstruction, 3D shape retrieval, object pose estimation, and scene layout optimization. We evaluate our system on both synthetic and real-world data to show we significantly outperform baselines from prior work. We also demonstrate generalization to internet images and the text-to-scene task.
Figure 1: We propose Diorama: a system for zero-shot single-view 3D scene modeling. Our system produces a holistic 3D scene model given a single input image, representing both architectural elements and objects in cluttered indoor scenes. We show three example applications of our system. The input image is shown in the top left in each case. Other images are renderings of the output 3D scene from different camera viewpoints. (a) 3D scene model from a synthetic image input. (b) 3D scene model from an internet image input. (c) Text-to-scene pipeline where the input image on the left is generated from a text prompt.
1 Introduction
Modeling a 3D scene from a single image input is a cornerstone task in computer vision that is critical in real-world applications such as autonomous driving, augmented/virtual reality, and robotic perception. Recent works tackling 3D perception as a reconstruction problem tend to produce imperfect surface-only 3D meshes that are incompatible with modern graphics pipelines or simulation platforms [64, 54, 9]. An alternative strategy leverages a database of CAD models to model a 3D scene in an analysis-by-synthesis manner [46, 31, 27] that enables convenient scene creation and editing. Unfortunately, the potential of this strategy is constrained by the limited availability of annotated image–CAD scene pairs and limited generalization to diverse real-world images.
Creating 3D scenes typically involves recruiting expert 3D designers to manually build and assemble the scenes. The resulting 3D scenes are often oversimplified [24] or of limited accessibility to the research community with only images available [104, 71]. More recent works address the synthetic data domain gap by aligning CAD annotations to real-world RGBD scans [4, 77, 55], RGB videos [60] or panoramas [89]. However, they either ignore architecture and scene structure or do not reach scales appropriate for training.
Recent studies show that foundation models exhibit 3D awareness [20, 100, 84]. This paper is motivated by the question “Is holistic 3D scene modeling from a single-view real-world image possible using foundation models?” To answer this question we present Diorama: a modular zero-shot open-world system that models synthetic 3D scenes given an image and requires no end-to-end training.
Our system consists of two major components: 1) Open-world perception for holistic scene understanding of the input image, including object recognition and localization, depth and normal estimation, architecture reconstruction and scene graph generation. 2) CAD-based scene modeling for assembly of a clean and compact 3D scene representation through CAD model retrieval, 9-DoF pose estimation, and semantic-aware scene layout optimization. Different from other zero-shot modular scene generation systems focusing on texture stylization [34], object reconstruction [106] and language description [26, 3], our work focuses on open-world holistic perception of objects, architecture, and spatial relationships. There are also works proposing to create synthetic data from images for robotic manipulation. However, current efforts [51, 13, 2, 86, 109, 10, 41] either use strong construction heuristics or only handle tabletop setups. We advocate our system as a more general approach to output cluttered compositional scenes that can be rearranged and edited. In summary, our contributions are:
  • We introduce Diorama, the first zero-shot open-world system for CAD retrieval-based 3D scene modeling from a monocular observation.
  • We ensure flexibility and generalization to real-world data with a modular design, proposing better and more robust solutions to general architecture reconstruction, 3D shape retrieval, zero-shot 9D pose estimation, and scene layout optimization.
  • We benchmark on synthetic and real-world data, and demonstrate the usability of our approach on in-the-wild images and text-to-scene applications.
Figure 2: Illustration of our Diorama system pipeline. The input image is processed in the open-world perception component (in orange box) through object instance segmentation, depth and normal estimation, architecture reconstruction and LLM-powered scene graph generation. The CAD-based scene modeling component (green box) then assembles a compositional 3D scene by retrieving and posing objects from a database and optimizing the overall scene layout. Multiple plausible scene arrangement hypotheses are produced as outputs.
2 Related work
We summarize prior work on open-world 3D perception, single-view scene modeling and reconstruction, and generative models for scene synthesis.
Open-world 3D perception. Visual foundation models enabled various aspects of open-world 3D perception, including 3D shape understanding [103, 108, 91, 92, 107, 50], part-level 3D shape recognition [56, 1], 3D semantic segmentation [67, 102, 8, 16, 39], 3D instance segmentation [78, 62, 36], 3D object detection [58, 96], neural field rendering [43, 68] and language grounding [40, 15]. Most works obtains semantic-rich representation for open-world language query by directly leveraging CLIP features or learning to align with CLIP features. Some works [32, 25] propose to learn tokenized 3D features along with LLMs. In contrast, our work directly models complex, cluttered 3D scene from a single-view image and unlike prior work does not rely on any 3D modality inputs such as point clouds.
Single-view scene modeling. There is a long history of work on scene modeling through CAD object retrieval and alignment [38, 35, 46, 47, 31, 48, 27, 59, 49]. IM2CAD [38] breaks down scene modeling into three major modules: room layout estimation, object recognition and scene configuration, similarly to our system. However, IM2CAD cannot easily generalize to unseen objects and oversimplifies complex architectures and arrangements in real-world images. More recent work explores end-to-end trained methods [46, 47, 31, 48]. Mask2CAD [46] jointly learns object detection, 3D shape embedding and direct pose regression. Patch2CAD [47] introduces a more robust patch-based joint embedding space for patch-level 3D shape retrieval. ROCA [31] leverages differentiable Procrustes optimization for pose estimation, and SPARC [48] employs an iterative “render-and-compare” strategy. DiffCAD [27] achieves state-of-the-art performance with a combination of probabilistic diffusion models trained in a weakly-supervised manner on synthetic data. However, all existing methods do not model architecture and require significant amounts of annotated data for training. Our system is training-free, requires no 3D supervision, and leverages large foundation models pretrained on large corpus internet data.
Single-view scene reconstruction. There is much work on single-view scene reconstruction [29, 37, 64, 54, 9, 101, 17]. Due to the limited availability of real-world data, prior works train on synthetic 3D shapes. This typically leads to limitations such as oversmoothed and incomplete meshes, discarding of fine-grained scene structure, and real to synthetic domain gaps. Recent work implements modular zero-shot scene reconstruction using inpainting and image-conditioned 3D shape generation for each object [106]. In contrast, we produce a holistic scene model including both architecture and objects, and fine-grained object relations in highly cluttered scenes not handled by this prior work.
Generative scene synthesis. Another line of work to synthetic 3D scene modeling is through learning scene configuration priors from large-scale data [24]. Various neural network-based approaches have explored capturing scene layout priors using CNNs [82, 70], relation graphs [83], transformers [66, 85], and diffusion models [87, 80, 52]. These approaches are restricted by training data quality and scale, and generally only produce furniture layouts on a horizontal plane. Recent work proposes to generate layout specifications using LLMs to mitigate the training data sparsity issues [26, 79, 28, 97, 3]. However, these approaches struggle to accurately position objects as LLMs do not encode fine-grained 3D spatial knowledge. In contrast, our system directly perceives images and converts them to plausible 3D layouts respecting the provided input.
3 System
We present Diorama, a zero-shot open-world system that reconstructs synthetic 3D scenes from monocular image inputs, without end-to-end training as required by prior work [31, 48, 27]. Our system has two major components: 1) Open-world perception for holistic scene understanding of the input image, including object recognition and localization, depth and normal estimation, architecture reconstruction and scene graph generation (Fig. 2, orange box). 2) CAD-based scene modeling for assembly of a clean and compact 3D scene representation through CAD model retrieval, 9-DoF pose estimation, and semantic-aware scene layout optimization (Fig. 2, green box). As a modular system, its performance can be naturally improved by replacing each individual component with a dedicated approach of better performance. We choose different components based on the principles of open vocabulary, category agnosticism, and robustness to non-exact object matches. The reconstructed scenes benefit from certain designs of Diorama, including scene graph for maintaining spatial relationships among objects, shape retrieval for multiple semantically similar arrangements, planar architecture for physically plausible supporting objects, and layout optimization for refining objects poses based on spatial relationships. We describe each module in the following sections, with additional details in the supplement.
3.1 Open-world Perception
Holistic scene parsing. Given an image $\mathbf{I}$, we recognize all objects in the image as a list of open-vocabulary categories $\mathbf{C}$. We combine an open-vocabulary detector OWLv2 [61] and Segment Anything [45] to localize object instance as bounding boxes $\mathbf{B}$ and masks $\mathbf{M}$. Thus, each localized object can be represented by its category, bounding box and instance mask, $\mathcal{O}=\{o_{i}=\{c_{i},b_{i},m_{i}\}\}_{i=1}^{N}$. We use Metric3DV2 [33] to estimate metric depth $\mathbf{D}$ and normal maps $\mathbf{N}$. We then back project the estimated depth to a point cloud $\mathbf{P}$. For each object $o_{i}$, we extract an instance point cloud $p_{i}$ from $\mathbf{P}$ using the bounding box $b_{i}$ and mask $m_{i}$.
[…]
Figure 3: Our PlainRecon architecture reconstruction approach. Objects are first segmented. The object masks are inpainted, and depth and normals are estimated. Then, normal-based clustering on the point clouds is used to produce the 3D architecture.
Planar architecture reconstruction. Reconstructing the surrounding architecture enables better physical plausibility of the scene arrangement, by ensuring objects are supported on surfaces and are not floating in the air. We represent architecture as 3D planes following [75, 72] to better capture general complex room layout in the real world. We introduce PlainRecon, a simple yet effective architecture reconstruction pipeline leveraging inpainting [99], monocular depth [33, 95] and normal estimation [33] that handles both synthetic and real-world images. PlainRecon consists of three steps (see Fig. 3): 1) segment objects and inpaint to produce an “empty room” showing the complete architecture; 2) extract a point cloud and normals for each architecture element using depth estimation; and 3) fit 3D planes to architecture elements using normal-based clustering. Given planar segments obtained from normals-based clustering, we fit a plane equation to each point cloud of an architectural element and compute bounds in each plane. See the supplement for more details.
Figure 4: Our zero-shot object pose estimation approach. We leverage vision transformer features to establish 2D and 3D correspondences and estimate 9-DoF poses for each object.
3.2 CAD-based scene modeling
Multimodal shape retrieval. We retrieve CAD objects from a database combining several 3D datasets [7, 11, 24, 14, 44]. Unlike 3D reconstruction pipelines, retrieval-based modeling creates a clean and compact scene representation which can more easily be used in downstream applications without post-processing, at the cost of potentially mismatches in geometry and appearance between 3D shapes and observed 2D objects. In essence, Diorama focuses on semantic matching of the observed objects under the open-world setting and focus on realistic object arrangements instead of precise geometric and texture modeling. We use DuoDuoCLIP [50] to encode text, images and 3D shapes into a joint embedding space. Specifically, we use a hierarchical retrieval strategy where we first retrieve from a text query to ensure that all 3D shape candidates lie in the same category group, then we sort retrieved shapes using an image query to re-rank shape candidates based on object appearance and take the best-matching one as the output.
Zero-shot object pose estimation. There is rich research literature in model-based pose estimation for unseen objects [30, 63, 88, 53, 5]. However, in our task we tackle a more challenging problem of estimating the pose of novel objects for semantically-matching but not necessarily exact geometrically-matching CAD objects. Our approach is a zero-shot category-level object pose estimation method similar to Goodwin et al. [30], that demonstrates generalization across unseen categories and removes the requirement for pose-labeled real image training data.
Specifically, we take advantage of the semantically-rich visual representation of a vision transformer [18] (ViT) learned in a self-supervised manner on large-scale data to compute object pose from feature similarity and correspondence. We pre-compute $T$ multiview renderings and depth maps for each retrieved CAD model $s_{i}$. Each query object crop $I^{o_{i}}$ and set of multiview images $\{I^{s_{i,j}}\}$ is encoded into normalized patch features $\mathcal{F}(I)$ using this ViT. For each patch embedding from the query image, we construct a semantic correspondence to a patch embedding from each rendering view with the minimal cyclical distance [30, 63]. We then compute a correspondence score as cosine similarity between patch embeddings and take the top-K correspondences. The query image is matched with the most similar multiview rendering based on the maximal averaged correspondence-wise feature similarity to produce a coarse pose hypothesis.
Since we have point clouds for both query object and CAD model we can also obtain 3D correspondences. Ideally, straightforward application of the Umeyama algorithm [81] with RANSAC on a set of 3D correspondences $q_{k}$ and $p_{k}$, can solve for a rigid body transform: rotation $R$, translation $t$, and a uniform scaling parameter $s$. However, we empirically find that $t$ and $s$ can be significantly influenced by undesired depth estimation and large object occlusion and truncation in image that result in visually implausible scene (object with extreme size or position deviation). We therefore adopt a small scale prediction network $\mathcal{F}_{\text{scale}}$ from GigaPose [63] for more robust scale estimation. The approach is illustrated in Fig. 4.
Figure 5: Our semantic-aware scene layout optimization. In this example a stack of books is placed on a side table.
Scene composition and optimization. We now have all the ingredients to compose a 3D scene from a set of retrieved and coarsely posed CAD models. Errors in the object poses may lead to violations in overall scene plausibility. Examples include inter-penetrations, flying objects, or gaps to the supporting objects. To address this, we use a differentiable semantic-aware scene layout optimization procedure to refine the coarse object poses. We separate the entire procedure into different stages to alleviate the difficulty of adjusting the 9D poses of all objects simultaneously. For each object, we first define its oriented bounding box (OBB) and then identify potential contact and support surfaces along with contact and support directions to optimize spatial support relations with other objects. We describe the optimization stages along with their terms below and illustrate the stage-wise optimization in Fig. 5.
Orientation. For a pair of objects $o_{i}$ and $o_{j}$, and their associated contact surface normal $\mathbf{n}_{i}^{c}$ and support surface normal $\mathbf{n}_{j}^{s}$, we define the term $e_{\text{align}}$ as the $L_{2}$ norm of difference between the two normals $e_{\text{align}}=\sum_{(i,j)\in G}||\mathbf{n}_{i}^{c}-\mathbf{n}_{j}^{s}||_{2}$. During pose refinement, we maintain the front direction of each object projected on the horizontal plane by penalizing large disturbances. Specifically, we define the term $e_{\text{sem}}$ as the $L_{2}$ norm of residuals added to the projected vector, $e_{\text{sem}}=\sum_{i\in V}||\mathbf{v}_{i}^{*}-\mathbf{v}_{i}||_{2}$ where $\mathbf{v}^{*}$ indicates the initial front vectors before optimization.
Placement. To ensure the supported object $o_{i}$ is placed on the supporting object $o_{j}$ for physical plausibility, we define the term $e_{\text{place}}$ as the $L_{1}$ norm of the distance vector from the center point of the contact surface to the support surface $e_{\text{place}}=\sum_{(i,j)\in G}||D_{\text{contact center}\rightarrow\text{support surface}}||_{1}$ We also define a spatial relation term $e_{\text{rel}}$ that deliberately maintains the relative $L_{2}$ distance between objects based on proximity relationships present in the scene, $e_{\text{rel}}=\sum_{(i,j)\in G}||(\mathbf{c}_{i}^{*}-\mathbf{c}_{j}^{*})-(\mathbf{c}_{i}-\mathbf{c}_{j})||_{2}$.
Space. Each object should occupy its own 3D space and not penetrate other objects. We measure the amount of collision among objects and architecture elements following Zhang et al. [101], using the Separating Axis Theorem (SAT) to calculate the collision between convex polygons (cuboids). We penalize the penetration depths between any two objects on three orthogonal axes, $e_{\text{col}}=\sum_{\{i,\;j\}\in V}\sum_{axis}d_{\text{penetration}}$. In particular, we handle the usually ignored but common case where objects are placed inside another object (e.g. books on a bookshelf) by defining a supporting volume for each object and an optimization term $e_{\text{vol}}$ that penalizes the summed protrusions of each object from its containing volume bounds.
Refinement. Lastly, we encourage all objects to stick to their supporting surface by running the placement stage again. Additionally, we establish an adherence relationship between each object on the floor and potential walls using a heuristic algorithm based on a distance threshold. We hence also optimize object position and rotation based on the adherence relationship.
4 Experiments
Figure 6: Examples on SSDB images. Scenes are modelled from estimated metric depth by Metric3Dv2 [33] and scene graphs are generated using GPT-4o. The third column visualizes the support hierarchy of the generated scene graphs. We predict object placement and arrangement in 3D. The “Arrangement 1” column uses ground-truth 3D shapes. “Arrangement 2” and “Arrangement 3” present similar semantic arrangements using different retrieved 3D shapes.
Ground Truth Depth Estimated Depth
Zero-shot pose estimation Scene-aware Alignment $\uparrow$ Collision $\downarrow$ Relation $\uparrow$ Scene-aware Alignment $\uparrow$ Collision $\downarrow$ Relation $\uparrow$
rAcc tAcc sAcc Acc rAcc tAcc sAcc Acc
Best-matching multiview 0.34 0.93 0.52 0.19 11.43 0.58 0.34 0.84 0.44 0.15 12.29 0.25
ZSP [30] w/ DinoV2 0.37 0.93 0.57 0.23 10.48 0.59 0.42 0.85 0.55 0.26 8.29 0.25
ZSP [30] w/ ft DinoV2 0.45 0.94 0.66 0.34 7.90 0.61 0.41 0.85 0.54 0.24 8.46 0.24
GigaPose [63] 0.36 0.95 0.71 0.27 7.91 0.61 0.36 0.84 0.64 0.22 7.44 0.22
Ours 0.45 0.95 0.71 0.33 7.26 0.61 0.41 0.84 0.64 0.25 7.37 0.25
Table 1: Comparison of zero-shot pose estimation methods for the 9-DoF CAD alignment task on SSDB images. GT objects are used to avoid 2D perception error for system analysis. We present results given GT and estimated depth.
Ablation Collision $\downarrow$ Scene Structure $\uparrow$
orientation placement overall
w/o optimization 7.26 0.18 0.01 0.00
+ Orientation 6.34 0.97 0.01 0.01
+ O, Placement 7.61 0.97 0.94 0.91
+ O, P, Space 5.90 0.97 0.54 0.53
+ O, P, S, Refinement 4.43 0.97 0.95 0.92
w/ GT scene graph 3.28 0.97 0.97 0.95
w/ GT architecture 4.15 0.97 0.97 0.95
Table 2: Ablation of the terms in our semantic-aware scene optimization. Scene graphs are generated with GPT-4o.
4.1 Experimental setup
Datasets. We evaluate on the Stanford Scene Database (SSDB) [23], which contains around 130 scenes. We choose this dataset as the scenes represent highly cluttered arrangements such as office desk setups and chemistry laboratories annotated using 1,700 unique CAD models ranging from large furniture (e.g. chair, table, bookshelf) to small items (e.g. handbag, clothing hanger, doll). We use Blender to render 344 scene images, with 3 camera views per scene on average. The SSDB dataset provides a rich support hierarchy annotation and defines the room architecture, allowing us to evaluate these two aspects in our output. To enlarge the 3D shape repository beyond what is present in SSDB, we include shapes from HSSD [44] and the LVIS set of Objaverse [14] to form an out-of-distribution (OOD) collection of 60K 3D shapes for retrieval.
Metrics. We evaluate the components of Diorama with a suite of metrics for object alignment, overall scene quality, and 3D shape retrieval. For alignment, we report the alignment accuracy for translation (tAcc), rotation (rAcc), and scale (sAcc), as well as overall accuracy (Acc). Prior work [31, 48, 27] judged a CAD model as correctly aligned if the translation error $\leq 20$cm, the geodesic rotation error $\leq 20^{\circ}$, and the scale ratio $\leq 20\%$. However, we found that scene plausibility can be severely reduced due to a few objects with unreasonable position, orientation or size even when average accuracy is high. Thus, we introduce a scene-aware alignment accuracy to penalize such cases. An object is considered correctly aligned under scene awareness if its relative position deviation $\leq 20^{\circ}$, relative rotation error $\leq 20^{\circ}$ and relative scale error $\leq 0.2$ to every other object in the scene. To measure the overall scene quality, we decompose the evaluation into two aspects: average mesh collision and scene structure hierarchy where we calculate the support relation prediction accuracy (where a prediction is correct if both support object and surface assignment are correct) and supportness accuracy where whether objects are properly positioned. To measure 3D shape retrieval similarity against the ground-truth shape, we compute the chamfer distance (CD) as in prior work [90, 27].
4.2 Quantitative Evaluation on SSDB
Models Household Furniture Occluded Complete Supported Supporting
SS OOD SS OOD SS OOD SS OOD SS OOD SS OOD
CLIP-H 8.5 11.2 5.1 8.8 8.5 11.3 7.3 10.2 8.2 11.1 6.0 8.9
OS-H 8.6 25.3 4.6 8.9 8.1 22.9 7.5 21.3 8.4 22.5 4.6 19.5
DD-T 9.4 12.4 5.9 10.9 9.1 12.6 8.3 11.6 9.2 12.3 6.5 10.8
DD-V 6.4 12.0 8.9 12.1 10.3 14.3 4.1 10.0 6.9 12.6 7.0 9.1
DD-H (Ours) 5.5 9.9 3.2 7.6 5.7 9.5 4.4 9.4 5.4 10.2 3.0 6.1
Table 3: Retrieval similarity (CD) results on SSDB images with top-5 retrievals, multiplied by $10^{3}$ for readability. We compare retrieving objects with different models - CLIP, OpenShape (OS), DuoduoCLIP (DD) - from SSDB (SS) and a wider set of out-of-distribution objects (OOD). We take the candidate with the minimum L1 Chamfer Distance to the ground truth. “T” and “V” mean using text or visual modality for query only respectively. “H” stands for using both modalities.
Figure 7: Qualitative results on ScanNet images. As detection, depth estimation and shape retrieval are modular components in both DiffCAD and Diorama, we use the same mask, depth and 3D shape inputs for alignment for fair comparison. We handle more categories in 1st and 4th examples (laptop and backpack).
Input Output Input Output
Figure 8: Examples for in-the-wild images. The output scenes capture large and small objects, and represent complex relations such as objects on walls (e.g., wall clock in top left) and objects on other objects (vase with flower on table, in the top right).
Text prompt Input Output Text prompt Input Output
Figure 9: Examples of applying Diorama in a text-to-scene setting. We generate an input image from a text prompt using Flux-1, and apply our system to the image. The output scenes conform to the text prompt.
We present evaluations of retrieval, object pose, and scene optimization. See the supplement for evaluation of the architecture reconstruction, and more ablations.
9D CAD alignment to image object. We compare our zero-shot pose estimation module with representative methods, including ZSP [30] and GigaPose [63], on the 9-DoF CAD alignment task on SSDB images. Since GigaPose originally cannot predict the size of CAD model, we augment it by integrating depth information. We include a baseline which only performs coarse pose selection using the best-matching multiview image with the highest average correspondence similarity. For ZSP, we use DinoV2 as the default feature extractor rather than Dino [6] for fair comparison. We also experiment with a variant of ZSP leveraging a fine-tuned DinoV2. Tab. 1 shows that our methods perform better with both ground-truth and estimated depth, as measured by metrics for alignment, collision and support relation prediction. ZSP performs slightly better on the overall alignment metric but significantly falls behind on the scale metric, which leads to implausible scene layouts in the end. The failure on scale estimation is due to the RANSAC-based pose solver that is substantially affected by noisy point correspondences.
Benefits of semantic-aware scene optimization. We conduct an ablation of the terms in the semantic-aware scene optimization. We set more strict thresholds for alignment accuracy where translation error $\leq$ 0.1cm and geodesic rotation error $\leq 1^{\circ}$. Tab. 2 shows that the full scene optimization (+ O,P,S,R) reduces collisions avoidance and remarkably improves scene hierarchical structure. The ablated results of gradually adding more optimization terms show that the orientation (O), placement (P), space (S) and refinement (R) terms all help improve object alignment. This table also reports results with a pseudo ground-truth scene graph and architecture to show that improvements to these modules can lead to better overall scene modeling.
3D shape retrieval results. Tab. 3 evaluates our multimodal 3D shape retrieval approach against state-of-the-art for open-world methods, including CLIP [69] and OpenShape [57]. We use the ViT-bigG-14 checkpoint of CLIP. We present results both for retrieving from a relatively small set of SSDB shapes, and from a large set of OOD shapes. Our approach outperforms others by a big margin reflecting that multiple modalities compensate for potential ambiguity in language or visual information loss in a photo of a complicated scene. The inferior performance of OpenShape using the same hybrid retrieval strategy implies that the multiview representation of 3D shapes we use is a better fit for the open-world problem setting.
4.3 Zero-shot Results on ScanNet
We validate our zero-shot open-world system on real-world images from ScanNet25k [12] and ShapeNet [7] CAD model annotations provided by Scan2CAD [4]. We deliberately evaluate on cluttered environments capturing occlusion and multiple real world objects. We do this by selecting 600 images from the standard evaluation set where each image contains at least 3 different object categories or 5 object instances. These images contain around 3,000 instances spanning 24 categories, compared to the evaluation set in prior work [42] with 2,500 instances from 2,100 images covering 6 categories. We compare to the state-of-the-art prior work ROCA [31] and DiffCAD [27] where ROCA is a fully-supervised deterministic end-to-end model trained on training data of Scan2CAD and DiffCAD is a weakly-supervised probabilistic model trained on generated synthetic data. However, we note that both these methods cannot scale up and generalize to unobserved real-world images as ROCA consumes expensive real-world data for training and DiffCAD requires training for each category individually and inefficient inference due to a denoising network design. In contrast, Diorama provides a training-free solution showing comparable performance against DiffCAD and more flexible capability against ROCA. We show a qualitative comparison of CAD alignment in Fig. 7. Note that we can handle out-of-distribution objects (laptop and backpack) in the first and fourth row, which ROCA and DiffCAD cannot process due to limited training data. See the supplement for quantitative results.
4.4 In-the-wild Results
Real-world image. We select a set of real-world images representing real-world object arrangement from pixabay, Pexels and Unsplash. We show qualitative results in Fig. 8.
Text-to-scene. We further expand the capability of Diorama to the text-to-scene problem. We first generate high-resolution realistic image given a text prompt using Flux-1 and apply our system to the image. We show qualitative results in Fig. 9.
5 Conclusion
Limitations. Open-world 3D perception through CAD model retrieval and alignment creates a compact 3D scene representation. However, applications requiring precise reconstructions may be hindered by the lack of exact geometry and texture matches. To address this, future work can investigate deforming and texturing the retrieved CAD models to improve alignment with observations while maintaining compactness, part-level structure and functionality. More advanced methods can be explored to learn affordance between objects for proper support. As RANSAC-based pose estimation can be inefficient and sensitive to noisy correspondences and large occlusions, stronger open-vocabulary pose estimation methods can be more preferable for real-time scene modeling applications.
We introduced Diorama, the first zero-shot open-world system that can holistically model a 3D scene from an RGB image without requiring end-to-end training or human annotations. Our system is composed of robust, generalizable solutions to each subtask of our problem: architecture reconstruction, 3D shape retrieval, 9D object pose estimation, and scene layout optimization. We evaluate our system on synthetic images and real-world data with quantitative analysis and visual comparisons, and demonstrate generalization to internet images and text-to-scene applications.
Acknowledgements. This work was funded in part by a CIFAR AI Chair, a Canada Research Chair, NSERC Discovery Grant, and NSF award #2016532. Daniel Ritchie is an advisor to Geopipe and owns equity in the company. Geopipe is a start-up that is developing 3D technology to build immersive virtual copies of the real world with applications in various fields, including games and architecture. We thank Xiaohao Sun, Jiayi Liu, Xingguang Yan, Chirag Vashist, and Wuyang Chen for feedback on the paper draft.
Appendix A Method details
For all experiments involving LMM-based visual reasoning, we use the checkpoint “gpt-4o-2024-08-06” of GPT-4o. It costs approximately \$0.12 on average per SSDB image for 344 examples.
A.1 Holistic scene parsing
Holistic scene parsing consists of several tasks to comprehensively understanding the semantics and geometry of a scene image that serve as inputs to downstream components, including object recognition and localization, and depth and normal estimation.
For open-world object recognition, given an image, we aim to identify the objects in the scene by prompting GPT-4o and run open-vocabulary detectors and segmentation models to obtain the bounding boxes and masks. Below, we provide prompts and additional details.
Prompt for identifying objects in an image: {minted}[breaklines, breaksymbol=, fontsize=]text What are objects in the image? Includes architectural elements such as floor and wall if applicable. Ensure that each class is singular and has no quantifiers.
Return the output in a JSON format according to the following format: ”classes”: [object1, object2, object3, …]
To obtain object bounding boxes and masks for localization, we first run the detector OWLv2 [61] by providing text inputs of a template prompt “a photo of CLASS” and obtain detection results after non-maximal suppression. To ensure each object instance is only captured by one detection box to avoid repeated 3D shape retrieval for the same instance, we apply class-aware multiple-instance suppression where the bounding box of the same category with the largest intersection-over-self (IoS) is discarded. We then prompt SAM [45] with each detected bounding box to predict the segmentation mask for each object by assigning the one with the highest score.
A.2 LMM-powered scene graph generation
Figure 10: Illustration of scene graph generation using GPT-4o.
Figure 11: We show the potential of Diorama in flexible scene editing in terms of different editing types, including object rearrangement, insertion, removal and replacement.
After obtaining the set of objects in the scene, we then use an LMM (e.g. GPT-4o) to obtain the scene-graph. The scene-graph provide information about relationships between objects, including support relations.
We specify to the LMM the desired output response in JSON format so it can be successfully executed by the program afterward. We also limit object supporting relations to be selected from [“placed on”, “mounted on”].
To encourage the LMM to obey the provided numeric object marks and reduce hallucination errors, we use interleaved text prompt by incorporating the object marks and template response format into the scene graph generation prompt directly for symbolic reference.
Prompt for extracting a scene graph for objects in a image. Note the objects were already identified and provided as input. {minted}[breaklines, breaksymbol=, fontsize=]text OBJECT_ENTRY = ”id”: obj_id (id of object), ”name”: obj_name (name of object), ”caption”: brief description of object appearance (excluding relationship), ”face_direction”: direction of the object (choose from the list face_direction), ,
SUPPORT_RELATION_ENTRY = ”subject_id”: obj_id (id of object), ”type”: choose from a list of supporting relationships support_rel, ”target_id”: id of supporting object, ,
Please follow the examples in the Visual Genome dataset and generate a scene graph that best describes the image. A reference list of labelled objects is provided as ”objects_with_ids” in the format of ¡object_id, object_name¿. Return the output in JSON according to the following format: ”room_type”: type of room, ”objects”: [ object_list ], ”support_relation”: [ support_relation_list ]
Relationships are expressed as ¡subject_id, type, target_id¿. Each pair of ¡subject_id, target_id¿ only occurs once.
A.3 Architecture reconstruction
To obtain a planar reconstruction of the architecture, we apply a three step process: 1) segmenting out objects and inpainting to obtain empty rooms, 2) using depth estimation to obtain 3D points, and 3) clustering points by normals and plane fitting to obtain planar architecture surfaces.
Segmentation and inpainting to obtain empty room. We begin by obtaining object masks that are used to inpaint the images, leaving us with unfurnished scenes. For SSDB, we use dichotomous segmentation methods such as BiRefNet [105] and MVANet [98]. To maximize the recall of predicted masks, we apply each method separately twice by filling the mask predicted in the first iteration with white color then re-running the models on such images and merging the masks from both iterations.
We find that BiRefNet excels at producing sharp and complete masks for large objects and dense object arrangements, while MVANet has higher recall for scenes without a single cluster of objects. Therefore, we opt to merge the masks produced by these two methods. Additionally, we use SEEM [110] for all other experiments as we find that scenes featuring objects heavily dispersed around the scene are challenging for dichotomous segmentation networks. We note that the benefit of using dichotomous segmentation compared to other segmentation networks is in completeness of the masks (e.g., they cover the full area of the object with higher probability). Ensuring that the masks are as complete as possible is crucial for the further steps of PlainRecon.
After obtaining a segmentation of the objects, we ‘erase’ the objects by applying inpainting to defurnish the room. We use LaMa [76] to inpaint the background (similar to Yu et al. [99]). While recent advances in generative models are producing higher-quality inpainting, we find that even specialized inpainting for removal methods such as CLIPAway  [19] still suffer from hallucinating rather than inpainting empty background. This step should ideally produce perfectly defurnished and sharp image. In practice, the resulting image is not fully defurnished and the inpainted region is blurry. While the former remains a bottleneck of the pipeline, we notice that the latter is frequently alleviated by robustness of monocular depth estimation models which is the next step of the pipeline.
Depth estimation to obtain point cloud. After we have a defurnished room, we apply depth estimation to obtain a 3D point cloud of the scene. We compare two depth estimation models applied on the inpainted images - DepthAnythingV2 (DAv2)  [95] and Metric3D (M3D)  [33]. As we require normals for clustering points into planes, we leverage M3D normal estimation capabilities. We note that the predicted normals are more robust to the blurriness of the inpainting step compared to depth. We project the pixels to point clouds using the predicted metric depth. At the end of this process, we have point clouds with normals of unfurnished scenes, where the variation in normals should now be exclusively explained by presence of different planes in scene architecture.
Plane identification and fitting. The next step is normals-based clustering of the points into planar segments. We begin by pre-processing point clouds by applying voxel downsampling and removing statistical outliers with both using implementations from Open3D [107]. Since voxel downsampling bins the points in each voxel into one, the resulting point cloud has an approximately even distance between pairs of neighboring points. This ensures stability of hyperparameters in the subsequent steps and filters some noise introduced by depth estimation models. Next we run K-means clustering ($k_{m}=12$ to account for noise) on normals to determine seed normals. We iteratively select the seed with the largest number of corresponding points, and cluster all the points that have normals within the angle threshold ($\alpha=10$ in our experiments). We further apply DBScan  [21] to separate the initial cluster as it is possible to have multiple walls with identical normals in the scene. The algorithm terminates when we either run out of seed normals or have less than the threshold number of unclustered points left ($N_{\min}=200$). Finally, we propagate the instance labels to full point clouds using KNN ($k_{n}=1$). We assign the $floor$ label to the largest cluster with a normal pointing sufficiently upwards. Similarly, we filter out ceiling clusters as those points with a downward facing normal.
Once we have obtained the point clusters, we fit bounded planar segments to the points to obtain the final architecture. We start by fitting a plane using RANSAC  [22]to obtain a plane equation, and then estimate the bounding boxes of architecture elements. To obtain tight boxes, we compute convex hulls of point cloud segments, project the convex hull vertices onto 2D plane and apply rotating calipers algorithm  [73] to obtain the bounding box with the smallest area. Finally, we convert vertices back into 3D. While such initialization provides good plane bounds, we need to refine the plane further to account for imperfections of previous steps. We begin by making sure that all the walls are orthogonal to the floor, this is simply done by solving a system of linear equations that finds a vector orthogonal to the intersection line of floor and wall and lies on the floor plane. We proceed to use this vector as a new normal for our wall equation. Then we proceed to refine the plane bounds by finding the intersections of each architecture element, projecting two closest vertices of one plane onto the intersection line and adjusting the vertices of the plane we projected onto to the projected vertices of the other plane, effectively connecting them. Finally, we duplicate the vertices with a small offset along the direction opposite to normal and export the meshes of architectural components.
Figure 12: Qualitative results on architecture reconstruction.
Method Depth # Success IoU $\uparrow$ PE $\downarrow$ EE $\downarrow$ RMSE $\downarrow$ CDb $\downarrow$ Time (s) $\downarrow$
RaC [75] DAv2 236 40.3 (58.8) 39.8 (12.2) 23.2 0.67 0.645 43
RaC [75] M3D 258 45.2 (60.3) 33.7 (11.6) 21.0 1.23 0.908 49
PlainRecon DAv2 344 47.8 13.3 27.5 1.26 0.503 23
PlainRecon M3D 344 58.6 9.6 18.9 1.37 0.447 29
Table 4: General architecture plane reconstruction results on SSDB. * and indicate estimated depth from Metric3DV2 [33] and DepthAnythingV2 [95] respectively. Numbers in parentheses indicate results of RaC when only consider successful predictions.
Groups #Cats / #Insts rAcc $\uparrow$ tAcc $\uparrow$ sAcc $\uparrow$ Acc $\uparrow$
Household 466 / 5577 0.44 0.84 0.71 0.29
Furniture 24 / 1137 0.53 0.84 0.69 0.34
Occluded 369 / 3175 0.43 0.82 0.64 0.25
Complete 401 / 3539 0.48 0.86 0.77 0.34
Supported 468 / 5715 0.46 0.85 0.71 0.30
Supporting 79 / 1634 0.46 0.82 0.70 0.28
Table 5: Averaged alignment results across different SSDB object groups. For each group, we count the number of fine-grained categories and object instances. Occluded objects are those with occlusion ratio above a threshold 5% of pixels.
Method supervision #hypo bed bkshlf cabinet chair sofa table bin bathtub display others cls. avg ist. avg
ROCA ✓✓ - 7.02 3.62 7.56 20.03 5.26 9.16 13.19 8.11 13.21 0.00* 8.72 12.87
DiffCAD 1 7.02 0.00 3.36 9.38 6.58 1.58 0.00* 0.00* 0.00* 0.00* 2.79 4.59
DiffCAD 5 12.28 0.72 6.72 14.2 7.89 1.58 0.00* 0.00* 0.00* 0.00* 4.34 6.84
DiffCAD 10 7.02 1.45 4.21 11.41 5.26 1.26 0.00* 0.00* 0.00* 0.00* 3.06 5.41
Ours - 0.00 2.96 3.39 8.62 3.95 3.63 0.00 0.00 8.21 4.76 3.55 5.87
Table 6: Object-focus alignment accuracy on the Scan2CAD benchmark. ROCA is fully-supervised using in-domain data. DiffCAD is weakly-supervised using synthetic data.
# retrievals Ground Truth Depth Estimated Depth
rAcc tAcc sAcc Acc rAcc tAcc sAcc Acc
1 0.21 0.93 0.46 0.09 0.21 0.82 0.42 0.08
4 0.32 0.96 0.64 0.18 0.32 0.85 0.59 0.15
8 0.37 0.96 0.69 0.21 0.36 0.84 0.65 0.18
Table 7: 9D CAD alignment results give a different number of retrieved 3D shapes using either ground-truth or estimated depth.
Ablation Collision $\downarrow$ Scene Structure $\uparrow$
orientation placement overall
all-in-one 12.49 0.18 0.84 0.15
stage-wise 4.43 0.97 0.95 0.92
Table 8: Comparison between all-in-one and stage-wise optimization.
Figure 13: Qualitative comparison between different ground-truth evaluation sets used in DiffCAD, ROCA and ours.
Figure 14: More examples on SSDB images.
Input Output Input Output
Figure 15: More examples for in-the-wild images.
Text prompt Input Output Text prompt Input Output
Figure 16: More examples of applying Diorama in a text-to-scene setting.
A.4 Multimodal 3D shape retrieval
For each observed object in the image, we use a two step process for retrieving matching shapes from our shape database. We first use a text query to retrieve objects matching the semantic category, and then image query to re-rank the retrieved candidates. We find this two-stage approach helps ensure that retrieved objects are of the correct semantic class, as it is possible for semantically different objects to be geometrically similar (e.g. books vs cardboard box).
We construct the text query from the template “a photo of CLASS” (similar to in detection phase). For the image query, we use object crop extracted from the original image and mask out background and occluders using bounding box $b_{i}$ and mask $m_{i}$. We also carefully separate retrieval of supporting and non-supporting objects according to the supporting hierarchy in the scene graph since supporting objects need to be pre-processed to represent only one object entity (not containing smaller sub-objects).
A.5 Zero-shot object pose estimation.
Given a pair of query image $I_{q}$ and reference rendering image $I_{r}$, we construct a “cyclical distance” map and define a similarity score using their corresponding DINOv2 patch features, $f_{q}$ and $f_{r}$, and feature masks $m_{q}$ and $m_{r}$. For each query feature $f_{q}^{i}$ at the patch location $i$, we compute its cyclical patch $i^{\prime}$ in $I_{q}$ as follows:
$\begin{split}i^{\prime}&=\operatorname*{argmax}_{w,m_{q}^{w}>0}S(f_{q}^{w},f_{r}^{j})\\ j&=\operatorname*{argmax}_{k,m_{r}^{k}>0}S(f_{q}^{i},f_{r}^{k})\end{split}$
where $S(\cdot,\cdot)$ denote as cosine similarity and a cyclical distance map is constructed as $D$ with $D_{i}=-||i,i^{\prime}||_{2}$. In the end, we build the feature correspondences by taking at most K query-reference patch pairs $(i,j)\in\mathcal{N}$ with top-K minimal cyclical distances in $D$. To ensure only keeping correspondences with strong similarity, we empirically set similarity score threshold as 0.7. The overall similarity between $I_{q}$ and $I_{r}$ is defined as the average of similarities of feature correspondences:
$\text{sim}(I_{q},I_{r})=\frac{1}{|\mathcal{N}|}\sum_{(i,j)\in\mathcal{N}}S(f_{q}^{i},f_{r}^{j})$
A.6 Scene optimization
During scene optimization, we take the retrieved objects, initial object pose estimates, and optimize the placements so that the relationships in the scene graph is respected. In the main paper, we described the different stages of optimization. Here we provide more details on how support is enforced while ensuring there is sufficient space, and no inter-penetrations.
Support. A key relationship is that of support, for which we want to ensure that the each object is properly place in contact with their supporting object. To do so, for each object in our database, we precompute candidate support surfaces by identifying roughly planar surfaces on the object.
Based on estimated initial object poses and the scene graph $\mathbf{G}$ describing object relationships, for each a pair of interacting objects, we compute a pair of contact surface from the supported object and support surface from the supporting object. The support surface is determined by selecting the surface with the minimal distance to the centre of the supported object. The contact surface is selected to be the one having the closest direction with the support direction.
Space. Particularly for the Space stage, we define a supporting volume for each object by extruding its identified support surface to an extent hitting another surface in the vertical direction. An object is properly supported only if it does not exceed the bounds of its corresponding supporting volume. We formulate the term $e_{\text{vol}}$ as the sum of distance from the corners of the contact surface to the sides of the supporting volume and the vertical distance between the centers of the object bounding box and the supporting volume.
Optimization. In each optimization stage, we use a separate SGD optimizer with initial learning rate 0.01 and momentum 0.9 for corresponding pose parameters, except for the Space stage where we set initial learning 0.001 for the scale parameter. We also decay the learning rate by 0.1 every 50 steps. We run 200 steps in total in each optimization stage.
Appendix B Additional experiments
Implementation details. We render SSDB scene images of size 1008$\times$784. Considering computation efficiency and good coarse pose selection, we render 180 gray-color multiviews of size 224$\times$224 for each 3D shape from predefined camera viewpoints to focus on geometry-wise semantic similarity and leave out effect of texture. For zero-shot pose estimation, we use the fine-tuned ViT-L of DinoV2 [65] following [63] to embed image patches of size $14\times 14$.
B.1 Architecture reconstruction evaluation
We compare our proposed PlainRecon against a recent method for obtaining 3D room layout via render-and-compare (RaC) [75]. RaC is a common architecture reconstruction baseline, and though follow-ups exist they introduce marginal improvements while having less reliable or no available public implementation [93, 72]. For a fair comparison, we provide RaC with inputs from more modern backbones compared to the ones used in the original implementation. We use DepthAnythingV2 or Metric3D depth and PlaneRecTR [74] planar segments.
Metrics. For architecture reconstruction, we report 2D IoU, Pixel Error (PE), Edge Error (EE) and Root Mean Squared Error (RMSE) following [75]. Additionally, we introduce bounding-box-centered chamfer distance (CDb) to measure the correctness of 3D structure of the predicted architecture. Additionally, we measure runtime for each method, averaged per scene.
Results. We compare our proposed PlainRecon with RaC in Tab. 4 on the 344 images of SSDB scene, with predicted depth from DepthAnythingV2 (DAv2) and Metric3D (M3D). We find that RaC fails in roughly 30% of the cases, while PlainRecon is able to handle all inputs. This is because RaC sets up architecture reconstruction as a constrained discrete optimization problem, which can occasionally produce infeasible solutions resulting in no output. PlainRecon does not involve such an optimization and therefore avoids this shortcoming. M3D depth is overall superior to DAv2 when used by any of the methods. PlainRecon closely matches or outperforms RaC on most of the metrics, RMSE being an outlier where PlainRecon underperforms significantly. Notably, PlainRecon outperforms in terms of CDb meaning that it captures the overall structure of the architecture better than RaC. As RaC fails in a significant number of scenes, PlainRecon is clearer superior in terms of reliability.
B.2 Performance for different object groups
We also analyze post-optimized object pose according to different object groups each object belongs to in Tab. 5. Since we aim for an open-world system that generalizes to long-tail categories in real life, we divide objects into three subgroups: household objects/common furniture, occluded/complete objects, and supported/supporting objects, rather than a coarse set of preselected categories as in prior work [31, 27]. We find that the performance is reduced for dominant household items, occluded objects, and supporting objects.
B.3 Multiple retrievals
In Tab. 7, we investigate the benefits of having more retrieved 3D shapes for correspondence computation and coarse pose proposal. It turns out that alignment accuracy increases with potentially more different pose initialization.
B.4 Different optimization strategy
We investigate the benefits of using a stage-wise optimization procedure rather than a more common all-in-one strategy where all terms are accumulated for optimizing simultaneously in Tab. 8. It shows that we obtain significant gains by decomposing the entire optimization task into separate stages.
B.5 Quantitative results on ScanNet
Tab. 6 shows quantitative comparison between ROCA, DiffCAD and ours on the proposed evaluation set. We note that ROCA is trained end-to-end using imperfect Scan2CAD annotations and DiffCAD is trained on the synthetic data per category. Both ROCA and DiffCAD cannot generalize to unseen objects during training (indicated as gray-colored numbers). Our zero-shot method achieves competitive performance against DiffCAD that further exhibits performance degradation under the probabilistic setting due to the partial observations of commonly occluded objects. Fig. 13 visually shows differences between ground-truth evaluation sets used in DiffCAD, ROCA and ours.
Appendix C Qualitative examples
We provide additional examples of generated scenes in Figures 14, 15 and 16. In Figure 14, we provide plausible arrangements based on renders from SSDB, as well as the associated predicted scene-graphs. With Diorama, we can produce alternative arrangments that uses different objects, while respecting the spatial relationships of the original image (e.g. picture on the wall, monitor on the desk).
We further showcase arrangements from real-world images (Figure 15) and images generated via text-to-scene (Figure 16).